find the ratius and center of the circle:
x^2+y^2-8x+14y+29=0

Answer by lwsshak3(11628)

find the ratius and center of the circle:
x^2+y^2-8x+14y+29=0
**
Standard form of equation for circle: (x-h)^2+(y-k)^2=r^2, (h,k) being coordinates of center and r=radius.
..
x^2+y^2-8x+14y+29=0
completing the square:
(x^2-8x+16)+(y^2+14y+49)=-29+16+49=36
(x-4)^2+(y+7)^2=36
Center: (4, -7)
Radius=√36=6